OpenStudy (anonymous):

Slope: Please explain to me what im doing wrong. I got the wrong answer. x=-7y-2. so I moved it around to -7y-2=x then added to to each side. +2 +2 and then I got -7y=x+2 divided each side by -7 -- -- -- -7 -7 -7 then I got y=-7-2/-7 But -7 isn't the slope :/

OpenStudy (johnweldon1993):

showed it in the message :)

OpenStudy (anonymous):

\[-7y=x+2\]\[\frac{ -7y }{ -7 }=\frac{ 1x }{ -7 }-\frac{ 2 }{ 7 }\]\[y=-\frac{ 1 }{ 7 }x-\frac{ 2 }{ 7 }\]

OpenStudy (anonymous):

-1/7 would be slope